# Does it matter? A more careful treatment of density fluctuations in 21-cm simulations

### v1

@@ IV Discussion
where $A_{v_{cb}}$ and $\beta_{v_{cb}}$ are free parameters, and $\langle v_{cb}^{2}\left(z_{\mathrm{rec}}\right)\rangle$ is the variance of the $v_{cb}$ field at recombination (see Eq. (16)). This modification in $M_{\mathrm{min}}$ results in having VAOs in the 21-cm power spectrum at large scales Muñoz (2019a, b); Sarkar and Kovetz (2023); Zhang et al. (2024). In addition to the above modification, 21cmFAST also modifies the current matter power spectrum by introducing a multiplicative correction factor of the form $1-A_{p}\exp\left[-\left(\ln\left[k/k_{p}\right]\right)^{2}/\left(2\sigma_{p}^{2}\right)\right]$ Muñoz et al. (2022). This factor arises from non-linear $v_{\mathrm{cb}}$ corrections, and was calibrated at $z=20$ with $A_{p}=0.24$, $k_{p}=300\,\mathrm{Mpc}^{-1}$, and $\sigma_{p}=0.9$. 21cmFirstCLASS has the option to remove this factor since it introduces non-linear corrections to $\delta_{m}^{\mathrm{lin}}$ and $\sigma\left(M_{h},z\right)$, while the EPS formalism requires these quantities to be evaluated from linear theory. Because this factor suppresses power at small scales, removing it advances cosmic dawn to slightly higher redshifts, as can be seen in Fig. 1 and 7. A second forewarning due here.
@@
Finally, a third nuance in the calculation arises due to the appearance of the smoothing radius $R$ in the RHS of Eq. (III.2). According to EPS, this $R$ should correspond to the region in which the local collapsed fraction is evaluated. Hence, we argue that the appropriate $R$ to consider should be associated with the simulation’s cell, $R=\left(3/4\pi\right)^{1/3}L_{\mathrm{cell}}$, where $L_{\mathrm{cell}}$ is the cell size of the low-resolution grid (for better precision, the realization of the perturbed box should be generated on a high-resolution grid). Note that the cell size cannot be arbitrarily small as we ought to have $M_{\mathrm{min}}<M_{R}$ for a non-zero local collapsed fraction (see square root in Eq. (III.2)). However, this condition is satisfied for a typical simulation cell of $R\sim L_{\mathrm{cell}}\sim 1\,\mathrm{Mpc}$; according to Eq. (22) such a smoothing radius corresponds to $M_{R}\sim 10^{11}\,M_{\odot}$ and even in ACGs (whose associated $M_{\mathrm{min}}$ is much larger than in MCGs) Barkana and Loeb (2001)
@@
V Conclusions
Building a robust theoretical framework for the 21-cm signal is challenging, especially during the cosmic dawn and reionization epochs, due to the involved complicated astrophysics of galaxy evolution and star formation. While former studies of the 21-cm signal have focused on refining the astrophysical models, they largely glossed over a trivial but crucial aspect of the signal: it is sourced from neutral hydrogen atoms in the IGM between recombination and reionization, and these hydrogen atoms are baryons. Therefore, even though baryons are coupled gravitationally to CDM at low redshifts, the relevant density field to consider is not the CDM density field, but rather the baryon density field. Approximating the baryon density field with the CDM density field when tracing the evolution of the neutral hydrogen atoms in the IGM, leads to misestimation of the 21-cm signal.
@@ V Conclusions
In this work, we distinguished between baryons and CDM to carefully study the cosmological impact of the baryons on the 21-cm signal. Unlike most studies which have adopted the SIGF approximation, whereby $\delta_{b}\propto D\left(z\right)$ and evolves in a scale-independent manner, we computed the baryon density field more precisely in 21cmFirstCLASS using the SDGF $\mathcal{D}_{b}\left(k,z\right)$. We have shown in previous work that this method agrees very well with Boltzmann-Einstein solvers at the dark ages when all perturbations can be evolved linearly. Here, we extended this method to the regime when the Universe entered its non-linear phase, including cosmic dawn when galaxies and stars begun to form, as well as reionization.
@@
We started our analysis from the dark ages, and demonstrated that the SIGF approximation fails poorly in that regime. We then moved to the non-linear epochs of cosmic dawn and reionization. While the exact physics in these epochs is still under uncertainty, especially as relates to the efficiency and production rate of stars and galaxies, two facts are certain regarding the cosmological role of the baryons in shaping the 21-cm signal: (1) their relative velocity with respect to CDM delays the signal, and (2) their non-linear clustering cannot be neglected at low redshifts and on small scales. In addition to understanding how the 21-cm signal depends on the baryon over-density $\delta_{b}$, we also discussed the role of the total matter over-density $\delta_{m}$ in the calculation. This enters through the collapsed fraction, Eq. (III.2), as part of the EPS formalism. While it is clear that the 21-cm signal at cosmic dawn and reionization should depend on the “non-linear” $\delta_{b}$ (i.e., $\delta_{b}$ as evaluated by non-linear theory), two common approaches are considered in the literature for the total matter density field $\delta_{m}$. One option is to use the “linear” $\delta_{m}^{\mathrm{lin}}$ since the EPS formalism is based on linear perturbation theory. Alternatively, one can choose to work with the non-linear $\delta_{m}^{\mathrm{NL}}$, as it was argued that this choice yields a better agreement with radiative transfer simulations. While Zeus21 (21cmFAST) adopts the former (latter) approach, 21cmFirtCLASS enables the user to decide which $\delta_{m}$ should be used in Eq. (III.2). Regardless, as was discussed in the previous section, all codes still make in their calculations some assumptions whose validity should be further examined in the future.
Efforts to measure the cosmological 21-cm signal are being carried out by various experiments. The expected data from these experiments, together with new photometric images from the James Webb Space Telescope Gardner et al. (2006); Mirocha and Furlanetto (2019); Harikane et al. (2023); Naidu et al. (2022); Castellano et al. (2022); Adams et al. (2023); Yan et al. (2023); Hassan et al. (2023), present an exciting opportunity to probe uncharted regions of our Universe, thereby deepening our understanding of the physical processes that governed its early evolution. Such an ambitious goal cannot be fulfilled without having robust theoretical predictions to be matched with the data. We thus believe that this work may serve to bring us one step closer towards this goal.

### v2

@@ II.5 Star formation rate density
where $A_{v_{cb}}$ and $\beta_{v_{cb}}$ are free parameters, and $\langle v_{cb}^{2}\left(z_{\mathrm{rec}}\right)\rangle$ is the variance of the $v_{cb}$ field at recombination (see Eq. 26). This modification in $M_{\mathrm{min}}$ results in having VAOs in the 21-cm power spectrum at large scales Muñoz (2019a, b); Sarkar and Kovetz (2023); Zhang et al. (2024), as was first studied in Refs. Dalal et al. (2010); Visbal et al. (2012); Fialkov et al. (2013).
III Methods
The main tool we use in this work to study how the 21-cm signal is sensitive to the assumed underlying matter fields is 21cmFirstCLASS. This code interfaces CLASS Blas et al. (2011) and 21cmFAST, allowing the simulation to begin from recombination with initial conditions that are consistent with the simulated cosmology, evolve the matter fields throughout the dark ages, and apply 2LPT during cosmic dawn and reionization to simulate the non-linear fluctuations of the signal in those epochs. Moreover, the merger between CLASS and 21cmFAST enables a consistent study of both CMB and 21-cm power spectra with 21cmFirstCLASS, even for cosmologies that extend beyond $\Lambda$CDM Flitter and Kovetz (2024a); Adi et al. (2024); Libanore et al. (2025).
@@ III Methods
For this work, three major modifications have been made in 21cmFirstCLASS with respect to its previous version, whose full description can be found in Ref. Flitter and Kovetz (2024a). Firstly, we enabled the option of using the linear matter density field $\delta_{m}^{\mathrm{lin}}$, rather the non-linear matter density field $\delta_{m}^{\mathrm{NL}}$ in the evaluation of the inhomogeneous SFRD $\dot{\rho}_{*}\left(\mathbf{x},z\right)$ (see discussion at the end of Sec. II.4). Secondly, we changed the effect that $v_{cb}$ has on the local SFRD; in addition to the modification of $M_{\mathrm{min}}^{\mathrm{MCG}}$ by $v_{cb}$ (Eq. 27), 21cmFAST also modifies the current matter power spectrum by introducing a multiplicative correction factor of the form $1-A_{p}\exp\left[-\left(\ln\left[k/k_{p}\right]\right)^{2}/\left(2\sigma_{p}^{2}\right)\right]$ Muñoz et al. (2022). This factor arises from non-linear $v_{\mathrm{cb}}$ corrections, and was calibrated at $z=20$ with $A_{p}=0.24$, $k_{p}=300\,\mathrm{Mpc}^{-1}$, and $\sigma_{p}=0.9$. 21cmFirstCLASS has the option to remove this factor since it introduces non-linear corrections to $\delta_{m}^{\mathrm{lin}}$ and $\sigma\left(M_{h},z\right)$, while the EPS formalism requires these quantities to be evaluated from linear theory, as we now do in 21cmFirstCLASS.
Figure 3: The 21-cm power spectrum as a function of redshift at $k=0.5\,\mathrm{Mpc}^{-1}$, when the baryon over-density field is evolved with the SIGF (blue curve) or the SDGF (orange curve). The inset shows a zoomed-in version of the curves after the dark ages, where a maximum difference of 6% can be seen.
Figure 4: Top left panel: realization of the baryon over-density field at $z=20$, evolved non-linearly with scale-dependent 2LPT. Top right panel: the fractional error in the normalized density field ($\Delta_{i}\equiv 1+\delta_{i}$) due to approximating $\delta_{b}\approx\delta_{c}$. The error is of the order of a percent at the most over- and under-dense regions. Bottom left (right) panel: realization of the baryon over-density field at $z=6$ when evolved linearly (non-linearly). Linear perturbation theory fails to capture the clustering of baryons at low redshifts and small scales.
The third modification that was made in 21cmFirstCLASS is relaxing the SIGF approximation in the calculations by accounting for the more exact SDGF. Assuming that linear perturbation theory holds during the dark ages, we have shown in Ref. Flitter and Kovetz (2024b) that implementing the SDGF in 21cmFirstCLASS would yield excellent agreement with fully relativistic Boltzmann-Einstein solvers such as CAMB, on sub-horizon scales that are larger than the Jeans scale (we verify the correctness of our calculations in Appendix D). On top of this modification during the dark ages, we improved the 2LPT calculations in 21cmFAST to account for scale-dependent growth also in the non-linear epochs of cosmic dawn and reionization. When relaxing the SIGF approximation, the displacement vector field depends on the considered matter species, and we denote it by $\mathbf{s}_{i}\left(\mathbf{q},z\right)$ ($i=b,c$). Now, the first order correction $\mathbf{s}_{i}^{(1)}\left(\mathbf{q},z\right)$ cannot be decomposed to spatial and temporal components, as in Eq. (15), and instead we must compute it in Fourier space,
When we relax the SIGF approximation in 21cmFirstCLASS, we use Eq. (28) for $\mathbf{s}_{i}^{(1)}$, instead of Eq. (15). On the other hand, the second order correction, $\mathbf{s}_{i}^{(2)}$, should be found by solving a set of coupled integro-differential equations, that the interested reader can find in Appendix C, Eq. (C.2). Instead, in 21cmFirstCLASS we keep using the approximation in Eq. (16); as described in detail in Appendix C, this is justified as by construction the second order correction $\mathbf{s}_{i}^{(2)}$ has a smaller contribution to $\mathbf{s}_{i}$, compared to the first order correction $\mathbf{s}_{i}^{(1)}$. We leave for future work the comparison of the approximated analytical solution of Eq. (16) with the fully numerical solution of Eq. (C.2).
IV Results
IV.1 Dark ages
During the dark ages, the only relevant matter species that contributes to the 21-cm signal are the baryons, as they source the emission and absorption of 21-cm radiation by hydrogen atoms, the recombination rate that governs $T_{k}$, and the coupling of $T_{s}$ to $T_{k}$ via collisional excitations. Thus the fluctuations in the 21-cm signal during the dark ages are entirely proportional to $\delta_{b}$. Because $\delta_{b}\ll 1$ during the dark ages, the scale-dependent evolution of the baryon density field barely affects the global 21-cm signal, i.e. the sky-averaged brightness temperature as a function of redshift, $\bar{T}_{21}\left(z\right)$. Indeed, for the cosmological and astrophysical parameters assumed in this work, we find that the SIGF approximation yields a maximum error of 0.4% in $\bar{T}_{21}$ at $z\simeq 14$, where the global signal reaches its minimum value.
where $P_{21}\left(k,z\right)$ is the angle-averaged Fourier transform of the two-point correlation function $\langle\delta_{21}\left(\mathbf{x},z\right)\delta_{21}\left(\mathbf{x}^{\prime},z\right)\rangle$, while $\delta_{21}$ is the local contrast in the brightness temperature, $\delta_{21}\left(\mathbf{x},z\right)\equiv T_{21}\left(\mathbf{x},z\right)/\bar{T}_{21}\left(z\right)-1$. In a previous work Flitter and Kovetz (2024b), we have shown that during the dark ages $\Delta_{21}^{2}\left(k,z\right)\propto\mathcal{T}_{b}^{2}\left(k,z\right)\ll\mathcal{T}_{c}^{2}\left(k,z\right)$. In Fig. 3, we show $\Delta_{21}^{2}\left(k,z\right)$, as calculated by 21cmFirstCLASS after the incorporation of the SDGF in the code. As expected, $\Delta_{21}^{2}\left(k,z\right)$ is suppressed during the dark ages when $\delta_{b}$ is correctly evolved with the SDGF. Specifically, we find that the SIGF approximation yields a $\sim 40\%$ error in the 21-cm power spectrum at $z=100$ for $k\gtrsim 0.1\,\mathrm{Mpc}^{-1}$.
IV.2 Cosmic dawn and reionization
As was previously discussed in Sec. II, when the Universe enters the cosmic dawn epoch, the dependence of the 21-cm signal on the matter fields becomes more complicated as (1) non-linear fluctuations in $\delta_{b}$ become non-negligible and (2) both $\delta_{b}$ and $\delta_{m}$ need to be taken into account.
We show in the top panels of Fig. 4 the comparison between the more precise $\delta_{b}^{\mathrm{NL}}$, computed with the scale-dependent first order LPT, and the common calculation that evolves the density field with the SIGF. At $z=20$, we see that the error is of the order of a percent at the most over- and under-dense regions. This error is expected to increase at higher redshifts due to the departure of the SDGF from the SIGF, while at lower redshifts it vanishes as $\mathcal{D}_{b}\left(k,z\right)\to D\left(z\right)$, see Fig. 1. The error in the evaluation of $\delta_{b}^{\mathrm{NL}}$ with the SIGF is enhanced when the 21-cm power spectrum is considered, due to the intricate interplay between non-linear fluctuations in different fields, e.g. $T_{k}$ and $J_{\alpha}$. Indeed, we see from comparing between the black and blue curves in Fig. 5 that the error in $\Delta_{21}^{2}$ may reach to $\mathcal{O}\left(10\%\right)$ at $z\gtrsim 12$. At low redshifts ($z\lesssim 10$), the error due to this approximation goes to zero because at these redshift $\delta_{b}$ approaches more closely to $\delta_{c}$.
The importance of evolving the baryon density field non-linearly is evident in the bottom panels of Fig. 4; while linear perturbation theory seems to give a rough estimate for the distribution of baryonic matter on very large scales, it fails to capture the clustering of baryons at low redshifts and small scales, and it wrongly predicts a more homogeneous density field. When computing the 21-cm power spectrum, the imprecise calculation from linear perturbation theory yields an order unity error at $z\lesssim 8$ at $k=0.3\,\mathrm{Mpc}^{-1}$, as demonstrated in Fig. 5 (this effect however is also manifest on other scales, as can be seen in Fig. 7). Importantly, this is the regime where upcoming 21-cm interferometers are mostly sensitive (c.f. Fig. 7).
Finally, Fig. 5 also demonstrates how $\Delta_{21}^{2}\left(k,z\right)$ is modified due to the choice of working with either $\delta_{m}^{\mathrm{lin}}$ or $\delta_{m}^{\mathrm{NL}}$ in Eq. (II.4). Evidently, because the fluctuations in the brightness temperature are strongly connected to the fluctuations in the SFRD, this choice has large implications on the resulting 21-cm power spectrum during cosmic dawn.
Figure 5: The 21-cm power spectrum as a function of redshift at $k=0.3\,\mathrm{Mpc}^{-1}$, when the baryon and the total matter over-density fields are evolved either linearly or non-linearly. The black curve corresponds to adopting the SIGF approximation in which $\delta_{b}=\delta_{c}=\delta_{m}$. The colors of the curves here match the same colors in Figs. 6-7, namely the blue, orange and green curves can be associated with 21cmFAST, Zeus21-like, and 21cmFirstCLASS, respectively. Here though, unlike Figs. 6-7, in all colorful curves we distinguish between baryons and CDM (i.e. in all colorful curves we use the SDGF), and we have the same $v_{cb}$ delaying effect (see more information at Sec. III).
Figure 7: Comparison of the 21-cm power spectrum on different scales, under the same conditions as in Fig. 6. Here, we show with error bars the expected noise from HERA in its design sensitivity at low redshifts ($z<10$) assuming the moderate foreground scenario, as evaluated by the 21cmSense code Pober et al. (2013, 2014) (we used the same 21cmSense settings as detailed in Ref. Flitter and Kovetz (2024a)). We also show a separate thick error bar at $z=14$, corresponding to a sensitivity estimate for SKA Meriot and Semelin (2024).
V Discussion
The results shown in the previous section demonstrate that differences in the simulated 21-cm signal can indeed arise by simply using different matter field configurations in the simulation. We now examine if these differences could be significant in future detections during cosmic dawn and reionization epochs. We thus compare between 21cmFirstCLASS and two popular codes used in the literature, 21cmFAST and Zeus21. Similarly to 21cmFirstCLASS, 21cmFAST performs 2LPT in order to compute the non-linear fluctuations in the matter fields, however there is no distinction between baryons and CDM and the code assumes $\delta_{b}=\delta_{c}=\delta_{m}$. Zeus21 also does not make the distinction between baryons and CDM, but unlike 21cmFAST it assumes the linear matter density field $\delta_{m}^{\mathrm{lin}}$ throughout the calculation. A summary of the differences between the codes is given in Table  1.
Feature 21cmFAST Zeus21 21cmFirstCLASS Diff. in $\Delta_{21}^{2}\left(k,z\right)$ Diff. in $\Delta_{21}^{2}\left(k,z\right)$
(cosmic dawn and reionization) (dark ages)
Non-linear evolution $\sim 20\%\left(z=15,k=0.3\,\mathrm{Mpc}^{-1}\right)$
Scale dependent growth $\sim 10\%\left(z=15,k=0.3\,\mathrm{Mpc}^{-1}\right)$ $\sim 40\%\left(z=100,k=0.3\,\mathrm{Mpc}^{-1}\right)$
Linear theory in EPS $\sim 20\%\left(z=7,k=0.1\,\mathrm{Mpc}^{-1}\right)$
Table 1: Summary of features studied in this work, comparing between three public codes: 21cmFAST, Zeus21 and 21cmFirstCLASS. The leftmost column shows the three features discussed in this work. “Non-linear evolution” corresponds to perturbing the total matter density field non-linearly, $\delta=\delta^{\mathrm{NL}}$. “Scale dependent growth” corresponds to the relaxation of the SIGF approximation, $\delta_{b}\neq\delta_{c}$. “Linear theory in EPS” corresponds to having $\delta_{m}=\delta_{m}^{\mathrm{lin}}$ in Eq. (II.4) while removing the non-linear $v_{cb}$ term in the calculation of $\sigma\left(M_{h},z\right)$ (see more details in Sec. III). The middle columns display whether or not these features are present in the three codes. The two rightmost columns show the relative difference in the 21-cm power spectrum due to switching off the features. During the dark ages, only the second feature affects the 21-cm power spectrum, thus 21cmFirstCLASS is the only code among the three that can simulate that epoch. The redshifts of our quoted benchmarks correspond to the minima of the global signal at $z=15$ and $z=100$, or the maximum at $z=7$. Before reionization ($z\gtrsim 7$), the relative differences do not vary significantly with scale between $0.1\leq k\leq 0.6\,\mathrm{Mpc}^{-1}$.
The comparison between the codes is shown in Figs. 6-7, where 21cmFAST (Zeus21) is represented with blue (orange) curve. As 21cmFirstCLASS is able to use either $\delta_{m}^{\mathrm{NL}}$ or $\delta_{m}^{\mathrm{lin}}$ (while always using the non-linear $\delta_{b}^{\mathrm{NL}}$ in the simulation after the dark ages are over), these cases are represented by the solid and dashed green curves, respectively. We note that in order to compare apples to apples, we obtained the orange “Zeus21-like” curve by running 21cmFAST with linear perturbation theory, which was shown to agree well with the Zeus21 code in Refs. Muñoz (2023); Cruz et al. (2024b).
In all three codes, the evaluated global signal is roughly the same, though minor differences can still be observed in Fig. 6. When 21cmFirstCLASS runs with the $\delta_{m}^{\mathrm{lin}}$ configuration, the signal is slightly shifted to higher redshifts, as $\sigma\left(M_{h},z\right)$ is calculated directly from linear perturbation theory, Eq. (20), without the non-linear $v_{cb}$ modification that suppresses power at small scales (see more details in Sec. III). Furthermore, the difference in the emission signal during the reionization epoch can get up to 12% when comparing 21cmFirstCLASS with $\delta_{m}^{\mathrm{lin}}$ to 21cmFAST (or to 21cmFirstCLASS with $\delta_{m}^{\mathrm{NL}}$). This discrepancy increases to 20% when Zeus21 is compared to 21cmFAST. These differences arise due to the highly non-linear fluctuations in that epoch; the fluctuations in the signal are so large that they do not cancel out when the mean is considered.
Next, in Fig. 7 we compare $\Delta_{21}^{2}\left(k,z\right)$, as evaluated by the three codes. In this figure, the $z\lesssim 10$ error bars indicate the expected noise level from HERA in its design sensitivity under the conservative assumption of a moderate foreground scenario, evaluated with 21cmSense Pober et al. (2013, 2014), while the $z=14$ error bar illustrates the sensitivity expected for SKA, estimated by Ref. Meriot and Semelin (2024). On large scales ($k\lesssim 0.3\,\mathrm{Mpc}^{-1}$), the 21-cm power spectrum features the same differences in the global signal at low redshifts, and the type of the assumed matter field ($\delta_{m}^{\mathrm{NL}}$ or $\delta_{m}^{\mathrm{lin}}$) is crucial for accurately simulating the signal within the noise level of HERA. This conclusion also holds for small scales ($k\gtrsim 0.3\,\mathrm{Mpc}^{-1}$), though to a lesser extent. Interestingly, at intermediate scales of $k\sim 0.3\,\mathrm{Mpc}^{-1}$ the type of the assumed matter field less matters, suggesting that at these scales the signal mostly depends on $\delta_{b}$ rather than $\delta_{m}$, though we comment that this conclusion might change if different astrophysical parameters are considered. In all scales, the usage of $\delta_{b}^{\mathrm{lin}}$ is inappropriate for simulating the signal within HERA sensitivity range, especially during the epoch of reionization. At higher redshifts, we see that distinguishing $\delta_{b}$ from $\delta_{c}$ causes differences in $\Delta_{21}^{2}\left(k,z\right)$ on large scales that are comparable to the expected SKA noise level.
@@ V Discussion
Fig. 7 thus demonstrates that the subtleties discussed in this work should matter in future analyses of 21-cm signal measurements. Yet, perhaps the most crucial subtlety in simulating the signal may come from another nuance which is treated the same in all three codes discussed here and has to do with the appearance of the smoothing radius $R$ in the RHS of Eq. (II.4). According to EPS, this $R$ should correspond to the region in which the local collapsed fraction is evaluated. Hence, we argue that the appropriate $R$ to consider should be associated with the simulation’s cell, $R=\left(3/4\pi\right)^{1/3}L_{\mathrm{cell}}$, where $L_{\mathrm{cell}}$ is the cell size of the low-resolution grid (for better precision, the realization of the perturbed box should be generated on a high-resolution grid). Note that the cell size cannot be arbitrarily small as we ought to have $M_{\mathrm{min}}<M_{R}$ for a non-zero local collapsed fraction (see square root in Eq. II.4). However, this condition is satisfied for a typical simulation cell of $R\sim L_{\mathrm{cell}}\sim 1\,\mathrm{Mpc}$; according to Eq. (24) such a smoothing radius corresponds to $M_{R}\sim 10^{11}\,M_{\odot}$ and even in ACGs (whose associated $M_{\mathrm{min}}$ is much larger than in MCGs) Barkana and Loeb (2001)
@@
VI Conclusions
Building a robust theoretical framework for the 21-cm signal is challenging, especially during the cosmic dawn and reionization epochs. Although valuable tools are available in the literature, all of them involve approximations or potential misinterpretations, often without thoroughly exploring their implications or consequences. For example, most of the codes model the 21-cm signal tracing the non-linear CDM density field, modeling its evolution using the SIGF. These choices glosses over a trivial but crucial aspect: since the 21-cm signal is sourced from neutral hydrogen atoms in the IGM, its modeling should follow the baryon density field. Even though baryons are coupled gravitationally to CDM at low redshifts, at higher redshifts this fact breaks down and the evolution of the two fields has to be carried on independently, also accounting for scale-dependent effects. Neglecting all of this potentially leads to misestimation of the 21-cm signal.
In this work, we compared two widely used public 21-cm codes, 21cmFAST Mesinger et al. (2011); Muñoz et al. (2022) and Zeus21 Muñoz (2023); Cruz et al. (2024b), with our code, 21cmFirstCLASS Flitter and Kovetz (2024a, b), an extension to 21cmFAST that enables to begin the simulation from recombination with any desired cosmology. We outlined the main approximations introduced by the former and relaxed them in our implementation, demonstrating their impact on the 21-cm global signal and power spectrum.
@@ VI Conclusions
We distinguished between baryons and CDM to carefully study the cosmological impact of the baryons on the 21-cm signal, similarly as in Ref. Naoz and Barkana (2005), but we also extended our analysis to lower redshifts, after galaxies were formed. Unlike most studies which have adopted the SIGF approximation, whereby $\delta_{b}\propto D\left(z\right)$ and evolves in a scale-independent manner, we computed the baryon density field more precisely in 21cmFirstCLASS using the SDGF $\mathcal{D}_{b}\left(k,z\right)$. We have shown in previous work that this method agrees very well with Boltzmann-Einstein solvers at the dark ages when all perturbations can be evolved linearly. Here, we extended this method to the regime when the Universe entered its non-linear phase, including cosmic dawn when galaxies and stars begun to form, as well as reionization.
We started our analysis from the dark ages, and demonstrated that the SIGF approximation fails poorly in that regime. We then moved to the non-linear epochs of cosmic dawn and reionization. As expected, even though some models of the SFRD consider the “linear” $\delta_{m}^{\mathrm{lin}}$ as an input to the conditional HMF, it is inconsistent to consider the “linear” $\delta_{b}^{\mathrm{lin}}$ once the dark ages are over, as this yields errors that are beyond the noise-level of 21-cm interferometers such as HERA. This is particularly true at low redshifts ($z\lesssim 8$) and large scales ($k\lesssim 0.3\,\mathrm{Mpc}^{-1}$), where the sensitivity of the interferometers is at maximum. Furthermore, we find that the SIGF approximation works very well for $z\lesssim 12$, as this is the regime where the baryons and CDM can indeed be characterized as a single fluid. In contrast, at higher redshifts, the SIGF approximation may lead to $\mathcal{O}\left(10\%\right)$ error, implying that advanced detectors like SKA will be sensitive enough such that the SIGF approximation is inadequate for simulating the signal in the required precision, even during cosmic dawn.
@@
In addition to understanding how the 21-cm signal depends on the baryon over-density $\delta_{b}$, we also discussed the role of the total matter over-density $\delta_{m}$ in the calculation. This enters through the collapsed fraction, Eq. (II.4), as part of the EPS formalism. While it is clear that the 21-cm signal at cosmic dawn and reionization should depend on the “non-linear” $\delta_{b}$ (i.e., $\delta_{b}$ as evaluated by non-linear theory), two common approaches are considered in the literature for the total matter density field $\delta_{m}$. One option is to use the “linear” $\delta_{m}^{\mathrm{lin}}$ since the EPS formalism is based on linear perturbation theory. Alternatively, one can choose to work with the non-linear $\delta_{m}^{\mathrm{NL}}$, as it was argued that this choice yields a better agreement with radiative transfer simulations. While Zeus21 (21cmFAST) adopts the former (latter) approach, 21cmFirtCLASS enables the user to decide which $\delta_{m}$ should be used in Eq. (II.4). Regardless, as was discussed in the previous section, all codes still make in their calculations some assumptions whose validity should be further examined in the future.
In conclusion, we can summarize our detailed comparison between the three codes as follows (see also Fig. 7 and Table 1). Zeus21 is a speedy code that attempts to reproduce the same 21-cm power spectrum as in 21cmFAST (albeit on much shorter timescales), but it only works with linear theory for the density field. While it is debatable whether or not it is more appropriate to consider the linear matter density field $\delta_{m}^{\mathrm{lin}}$ in the EPS and SFRD calculations, it is clear that it is inconsistent to consider the linear baryon density field $\delta_{b}^{\mathrm{lin}}$ at low redshifts, where HERA and other 21-cm interferometers are most sensitive. Even though the exact separation between baryons and CDM is not critical for analyzing the data from current interferometers, it would become increasingly more important in the future when data from next generation interferometers like SKA becomes available. At the moment, 21cmFirstCLASS is the only semi-numerical code that treats separately and consistently the different evolution in the baryon and CDM density fields during cosmic dawn and reionization epochs. Since 21cmFirstCLASS is also capable of simulating the dark ages, where this effect is most pronounced, it is thus poised to serve as an important tool in analyzing the 21-cm signal as measured by future interferometers, whether they are ground-based or lunar-based.